A jar contains four red, four yellow, and three green jelly beans. If Joan and Jim take one jelly bean each, the probability that they bot take a red jelly bean is?
Thank you for your help

Answer by robertb(4012)

(4/11)*(3/10) = 12/110 = 0.1091, to 4 decimal places.